# Accelerating Model Predictive Control for Legged Robots through Distributed Optimization

### v3

@@ Abstract
This paper presents a novel approach to enhance Model Predictive Control (MPC) for legged robots through Distributed Optimization. Our method focuses on decomposing the robot dynamics into smaller, parallelizable subsystems, and utilizing the Alternating Direction Method of Multipliers (ADMM) to ensure consensus among them. Each subsystem is managed by its own Optimal Control Problem, with ADMM facilitating consistency between their optimizations. This approach not only decreases the computational time but also allows for effective scaling with more complex robot configurations, facilitating the integration of additional subsystems such as articulated arms on a quadruped robot. We demonstrate, through numerical evaluations, the convergence of our approach on two systems with increasing complexity. In addition, we showcase that our approach converges towards the same solution when compared to a state-of-the-art centralized whole-body MPC implementation. Moreover, we quantitatively compare the computational efficiency of our method to the centralized approach, revealing up to a 75% reduction in computational time. Overall, our approach offers a promising avenue for accelerating MPC solutions for legged robots, paving the way for more effective utilization of the computational performance of modern hardware. Accompainying video at https://www.youtube.com/watch?v=0KcTnGYjJPw
@@ I Introduction
Planning and control for legged systems presents significant challenges due to their inherently nonlinear dynamics. Designing trajectories for such systems is a demanding task, as these trajectories must satisfy system constraints, including the robot’s kinematic limits, actuator torque limits, and the constraints imposed by the surrounding environment, such as obstacles or gaps. Model Predictive Control (MPC) [1] has emerged as a powerful technique for generating complex motions in robotic systems, particularly in the domain of legged robots [2]. MPC approaches calculate the optimal control inputs based on the modeled system, resulting in complex constrained optimization problems that need to be solved rapidly enough to achieve online re-planning of the optimal trajectory. Recent advancements in onboard hardware computational power and the development of more efficient Nonlinear Programming (NLP) solvers like [3],[4], and [5] have opened doors to solving efficiently complex Optimal Control Problem (OCP). However, the solution time is still a crucial factor to be considered when designing a control architecture for legged systems. To mitigate the computational complexity associated with the whole-body dynamics of the robot and speed up the solving time of the OCP, the works [6], [7], and [8] opted to approximate the robot with the centroidal model [9]. This approach reduces the robot’s dynamics to its center of mass (CoM) motion, neglecting the influence of individual limbs on the base. Although this approximation may seem simplistic, it considers the key dynamic effects on the system. Additionally, it facilitates the integration of feasibility constraints, such as contact friction. The centroidal model has demonstrated effectiveness in stabilizing locomotion for both quadrupeds and bipeds, although with the necessity for carefully crafted reference trajectories to track agile motions.
@@
To reduce the drawbacks of such a model, [10] and [11] developed an MPC formulation based on the Kino-Dynamic model, which takes into account joint position and velocity, neglects the effect of limb acceleration on the base dynamics while maintaining the effect of the robot joint configuration on the inertia matrix. As a result, the Kino-Dynamic model avoids the highly non-linear terms while simultaneously providing a better approximation than the centroidal model. In contrast, [12] and [13] used the full-body model, entirely exploiting the robot’s capabilities. The complexity of the OCP s considered in those implementations is remarkable given the limited time budget available for solving them in the feedback MPC scheme. Indeed, it needs to be considered that the computational complexity of the NLP implementation increases with the number of state variables in the OCP, typically following a cubic $O(n^{3})$ law [14]. This poses a significant limitation on the number of variables that can be feasibly included in the optimization process and restricts the length of the MPC horizon. Despite the continuous growth of the number of CPU cores of each generation of modern processors, the algorithms presented earlier face limitations in terms of parallelizability. This limits the effective utilization of the computational resources and restricts how the solver’s solution time scales with the number of decision variables. A different paradigm that tries to reduce the computation burden is presented in [15]. The authors worked towards a reduction of the computational effort by decoupling a quadruped robot into two bipeds. This brought a consistent reduction in the number of optimization variables that drastically decreased the computational time; nevertheless, the formulation they employed limits the approach to only the offline generation of periodic orbits, too slow to be computed online in a feedback MPC scheme. In contrast, [16] proposed a similar decoupling principle and applied it to quadratic programming-based nonlinear controllers. While the parallelizability of their approach leads to remarkable performances, particularly in terms of disturbance rejection properties inherited from the decoupling framework, their implementation requires the knowledge of the centralized system optimum. The approach is thus limited to tracking pre-computed periodic orbits. Finally, [17] developed local controllers based on the dynamic decomposition from [15], which can stabilize the centralized system across various scenarios. However, their approach is not suitable for generating highly dynamic motions due to the intrinsic limitations of instantaneous controllers. The above-mentioned methods also lack easy adaptability to more complex systems, such as a quadruped robot with one or several articulated arms on top.
@@
Our approach divides the robot into multiple subsystems while ensuring coherence between the solutions of the decoupled systems with a consensus formulation. In particular, we designed an easily scalable method to speed up the computation of the OCP on legged systems exploiting a parallelized implementation of the Alternating Direction Method of Multipliers (ADMM). It is worth highlighting that ADMM has already been successfully used in legged robotics by [18]. Their work primarily focused on rigorously enforcing consensus between simplified and whole-body optimizations within a two-step optimization framework, which is common in locomotion controllers. In contrast, our approach focuses on decomposing the robot dynamics into reduced subsystems and using ADMM to ensure consensus among them. In other words, our approach boils down to running in parallel a separate MPC for each sub-system we have divided our robot into, using consensus ADMM to maintain consistency between the optimizations. This allows our algorithm to scale effectively with more complex robot configurations. For instance, the issue of integrating an articulated arm onto a quadruped is simplified to merely adding another subsystem in parallel to the others and adding the relative consensus. By adopting this methodology, our method is able to solve complex whole-body motions (Fig. 1), overcoming at the same time the curse of dimensionality that is normally associated with the use of the robot’s full dynamics.
@@ I-A Contributions
  • a quantitative evaluation in simulation of the performance and convergence property of the proposed method against a state-of-the-art implementation of the same whole-body MPC problem.
@@ II-B Model Predictive Control
$\begin{split}&\boldsymbol{v}_{k+1}=\boldsymbol{v}_{k}-\boldsymbol{M}^{-1}(\boldsymbol{b}-\boldsymbol{S}\boldsymbol{\tau}-\boldsymbol{J}^{T}\boldsymbol{\lambda})dt\\ &\boldsymbol{q}_{k+1}=\boldsymbol{q}_{k}\oplus\boldsymbol{v}_{k+1}\end{split}$ (5)
@@
$\min\limits_{\boldsymbol{x},\boldsymbol{u}}\quad$ $\boldsymbol{l}_{T}(\boldsymbol{x}_{N}+\sum_{k=0}^{N-1}{\boldsymbol{l}(\boldsymbol{x}_{k},\boldsymbol{u}_{k})}$ (6a)
@@
$\boldsymbol{q}_{k+1}=\boldsymbol{q}_{k}\oplus\boldsymbol{v}_{k+1}$ (6c)

### v5

@@ Abstract
This paper presents a novel approach to enhance Model Predictive Control (MPC) for legged robots through Distributed Optimization. Our method focuses on decomposing the robot dynamics into smaller, parallelizable subsystems, and utilizing the Alternating Direction Method of Multipliers (ADMM) to ensure consensus among them. Each subsystem is managed by its own Optimal Control Problem, with ADMM facilitating consistency between their optimizations. This approach not only decreases the computational time but also allows for effective scaling with more complex robot configurations, facilitating the integration of additional subsystems such as articulated arms on a quadruped robot. We demonstrate, through numerical evaluations, the convergence of our approach on two systems with increasing complexity. In addition, we showcase that our approach converges towards the same solution when compared to a state-of-the-art centralized whole-body MPC implementation. Moreover, we quantitatively compare the computational efficiency of our method to the centralized approach, revealing up to a 75% reduction in computational time. Overall, our approach offers a promising avenue for accelerating MPC solutions for legged robots, paving the way for more effective utilization of the computational performance of modern hardware. Accompanying video at https://youtu.be/Yar4W-Vlh2A. The related code can be found at https://github.com/iit-DLSLab/DWMPC
@@ I Introduction
Planning and control for legged systems presents significant challenges due to their inherently nonlinear dynamics. Designing trajectories for such systems is a demanding task, as these trajectories must satisfy system constraints, including the robot’s kinematic limits, actuator torque limits, and the constraints imposed by the surrounding environment, such as obstacles or gaps. Model Predictive Control (MPC) [1] has emerged as a powerful technique for generating complex motions in robotic systems, particularly in the domain of legged robots [2]. MPC approaches calculate the optimal control inputs based on the modeled system, resulting in complex constrained optimization problems that need to be solved rapidly enough to achieve online re-planning of the optimal trajectory. Recent advancements in onboard hardware computational power and the development of more efficient Nonlinear Programming (NLP) solvers like [3],[4], and [5] have opened doors to solving efficiently complex Optimal Control Problem (OCP). However, the solution time is still a crucial factor to be considered when designing a control architecture for legged systems. To mitigate the computational complexity associated with the whole-body dynamics of the robot and speed up the solving time of the OCP, the works [6], [7], and [8] opted to approximate the robot with the centroidal model [9]. This approach reduces the robot’s dynamics to its center of mass (CoM) motion, neglecting the influence of individual limbs on the base. Although this approximation may seem simplistic, it considers the key dynamic effects on the system. Additionally, it facilitates the integration of feasibility constraints, such as contact friction. The centroidal model has demonstrated effectiveness in stabilizing locomotion for both quadrupeds and bipeds, although it is necessary to carefully craft reference trajectories to track agile motions.
@@
To reduce the drawbacks of such a model, [10] and [11] developed an MPC formulation based on the Kino-Dynamic model, which takes into account joint position and velocity, neglects the effect of limb acceleration on the base dynamics while maintaining the effect of the robot joint configuration on the inertia matrix. As a result, the Kino-Dynamic model avoids the highly non-linear terms while simultaneously providing a better approximation than the centroidal model. In contrast, [12] and [13] used the full-body model, entirely exploiting the robot’s capabilities. The complexity of the OCP s considered in those implementations is remarkable, given the limited time budget available for solving them in the feedback MPC scheme. Indeed, it needs to be considered that the computational complexity of the NLP implementation increases with the number of state variables in the OCP, typically following a cubic $O(n^{3})$ law [14]. This poses a significant limitation on the number of variables that can be feasibly included in the optimization process and restricts the length of the MPC horizon. Despite the continuous growth of the number of CPU cores of each generation of modern processors, the algorithms presented earlier face limitations in terms of parallelizability. This limits the effective utilization of the computational resources and restricts how the solver’s solution time scales with the number of decision variables. A different paradigm that tries to reduce the computation burden is presented in [15]. The authors worked towards a reduction of the computational effort by decoupling a quadruped robot into two bipeds. This brought a consistent reduction in the number of optimization variables that drastically decreased the computational time; nevertheless, the formulation they employed limits the approach to only the offline generation of periodic orbits, too slow to be computed online in a feedback MPC scheme. In contrast, [16] proposed a similar decoupling principle and applied it to quadratic programming-based nonlinear controllers. While the parallelizability of their approach leads to remarkable performances, particularly in terms of disturbance rejection properties inherited from the decoupling framework, their implementation requires the knowledge of the centralized system optimum. The approach is thus limited to tracking pre-computed periodic orbits. Finally, [17] developed local controllers based on the dynamic decomposition from [15], which can stabilize the centralized system across various scenarios. However, their approach is not suitable for generating highly dynamic motions due to the intrinsic limitations of instantaneous controllers. The above-mentioned methods also lack easy adaptability to more complex systems, such as a quadruped robot with one or several articulated arms on top.
@@
Our approach divides the robot into multiple subsystems while ensuring coherence between the solutions of the decoupled systems with a consensus formulation. In particular, we designed an easily scalable method to speed up the computation of the OCP on legged systems exploiting a parallelized implementation of the Alternating Direction Method of Multipliers (ADMM). It is worth highlighting that ADMM has already been successfully used in legged robotics by [18]. Their work primarily focused on rigorously enforcing consensus between simplified and whole-body optimizations within a two-step optimization framework, which is common in locomotion controllers. In contrast, our approach focuses on decomposing the robot dynamics into reduced subsystems and using ADMM to ensure consensus among them. In other words, our approach boils down to running in parallel a separate MPC for each subsystem we have divided our robot into, using consensus ADMM to maintain consistency between the optimizations. This allows our algorithm to scale effectively with more complex robot configurations. For instance, the issue of integrating an articulated arm onto a quadruped is simplified to merely adding another subsystem in parallel to the others and adding the relative consensus. By adopting this methodology, our method is able to solve complex whole-body motions (Fig. 1), overcoming at the same time the curse of dimensionality that is normally associated with the use of the robot’s full dynamics.
@@ I-A Contributions
  • a quantitative evaluation, in simulation of the performance and convergence property of the proposed method against a state-of-the-art implementation of the same whole-body MPC problem.
@@ II-B Model Predictive Control
$\begin{split}&\boldsymbol{v}_{k+1}=\boldsymbol{v}_{k}-\boldsymbol{M}^{-1}(\boldsymbol{b}-\boldsymbol{S}\boldsymbol{\tau}-\boldsymbol{J}^{T}\boldsymbol{\lambda})dt\\ &\boldsymbol{q}_{k+1}=\boldsymbol{q}_{k}\oplus\boldsymbol{v}_{k+1}dt\end{split}$ (5)
@@
$\min\limits_{\boldsymbol{x},\boldsymbol{u}}\quad$ $\boldsymbol{l}_{T}(\boldsymbol{x}_{N})+\sum_{k=0}^{N-1}{\boldsymbol{l}(\boldsymbol{x}_{k},\boldsymbol{u}_{k})}$ (6a)
@@
$\boldsymbol{q}_{k+1}=\boldsymbol{q}_{k}\oplus\boldsymbol{v}_{k+1}dt$ (6c)